OpenStudy (anonymous):

Write in vertex form: y=x^2 -4x-1

OpenStudy (helder_edwin):

complete the square.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i know what im supposed to do i dont know how to do it...

OpenStudy (helder_edwin):

\[ \large y=x^2-4x-1 \] \[ \large y+1\color{red}{+4}=x^2-4x\color{red}{+4} \] \[ \large y+5=(x-2)^2 \]

OpenStudy (anonymous):

oh thank you so so much!!

OpenStudy (helder_edwin):

u r welcome did u get it?

OpenStudy (anonymous):

were did the 1 go? and how did you get the 4?

OpenStudy (helder_edwin):

the 1 is negative first. when it changes to the left side it becomes positive

OpenStudy (anonymous):

okay. i think i get it thank you

OpenStudy (helder_edwin):

the 4 (in red) is the square of the half of the coefficiente of x (which is -4).

OpenStudy (anonymous):

oh okay yeah i get it now thank you